Given a hypothetical consumption function of the form:
Y = C + I0 + G0
C = α + β
Where: = Y – T
Y = Income
T = Taxes and that:
Compute the equilibrium level of income and consumption
1
Expert's answer
2016-11-24T12:14:10-0500
Solution:
Y = α + β + (Y – T) + G0
Y = C
α + β + (Y – T) + G0 = α + β
Y = C = T – G0